Calculate the density of a mixture of methanol (CH3OH) and water at 20oC having a
composition of 40% w/w methanol (State your assumption)
1
Expert's answer
2020-06-25T07:51:55-0400

The density of any substance and mixture can be calculated using the following formula:

d = m / V

where d - density, m - mass, V - volume

Let assume that the volune if the mixture is 100 ml.As the temperature is 20 °C, the density of pure water is 1 g/ml, and density of pure methanol is 0.791 g/ml.

From here the mass of pure water and methanol in 100 ml equals:

m(water) = d(water) × V(water) = 1 g/ml × 100 ml = 100 g.

m(methanol) = d(methanol) × V(methanol) = 0.791 g/mol × 100 ml = 79.1 g.

As the mixture contains 40% methanol and 60 %, it consists of 0.40 × 79.1 = 31.64 g methanol and 0.60 × 100 g = 60 g water.

From here, the total mass of the mixture is:

m(mixture) = 32.64 g + 60 g = 92.64 g

Finally, the density of the mixture is:

d(mixture) = 92.64 g / 100 ml = 0.93 g/ml


Answer: 0.93 g/ml